# Polynomial stability of non-linearly damped contraction semigroups

### v1

@@ 3.1. The wave equation with weak non-linear damping
to be solved subject to the boundary conditions $u(0,t)=u(1,t)=0$ for all $t>0$ and the initial conditions $u(\cdot,0)\in H_{0}^{1}(0,1)$, $u_{t}(\cdot,0)\in L^{2}(0,1)$. Here the function $b\in L^{2}(0,1;\mathbb{R})$ models the presence of weak (distributed) damping and the function $\phi:\mathbb{C}\to\mathbb{C}$ is potentially non-linear. We may formulate the problem in the form of (1.1) for the state variable $x(t)=(u(\cdot,t),u_{t}(\cdot,t))$, $t\geq 0$, by setting $X=H_{0}^{1}(0,1)\times L^{2}(0,1)$, $U=\mathbb{C}$, choosing $A$ to be the operator defined by $A(u,v)=(v,u^{\prime\prime})$ for all $(u,v)$ in the domain $D(A)=(H^{2}(0,1)\cap H_{0}^{1}(0,1))\times H_{0}^{1}(0,1)$ and defining $B\in\mathcal{B}(U,X)$ by $Bz=(0,b(\cdot)z)$ for all $z\in\mathbb{C}$. We denote the (rescaled) Fourier sine series coefficients of $b$ by
@@
Consider the system (1.1) for the weakly damped wave equation as above, and suppose there exists $\beta>0$ such that $|b_{n}|\gtrsim n^{-\beta}$ for all $n\geq 1$. Suppose further that $\phi:U\to U$ is monotone and locally Lipschitz continuous, and that (1.2) holds for some $\delta>0$. Then all generalised solutions of (1.1) satisfy $\|x(t)\|\to 0$ as $t\to\infty$, and all classical solutions satisfy $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$.
@@
The operator AA is skew-adjoint, and its spectrum consists of the simple eigenvalues $i\pi n$ for $n\in\mathbb{Z}\setminus\{0\}$, which are uniformly separated. As in the proof of [11, Cor. 6.3], our assumption on the decay of the Fourier sine series coefficients of $b$ implies that $i\mathbb{R}\subseteq\rho(A_{B})$ and $\|(is-A_{B})^{-1}\|\lesssim 1+|s|^{2\beta}$ for all $s\in\mathbb{R}$. The result now follows from Corollary 2.6. ∎
@@
    Note that in the particular case where $\phi$ is the identity map on $\mathbb{C}$, so that (1.2) holds for $\delta=1$, Proposition 3.1 is sharp in the sense that if $\limsup_{n\to\infty}n^{\beta}|b_{n}|>0$ then for any function $r:\mathbb{R}_{+}\to(0,\infty)$ such that $r(t)=o(t^{-1/(2\beta)})$ as $t\to\infty$ there exists a classical solution such that $\|x(t)\|\neq O(r(t))$ as $t\to\infty$. Indeed, if this is not the case then an application of the uniform boundedness principle gives $\|T_{B}(t)\smash{A_{B}^{-1}}\|=O(r(t))$ as $t\to\infty$. On the other hand, since $\limsup_{n\to\infty}n^{\beta}|b_{n}|>0$ it follows from [11, Prop. 5.1] that
    $\limsup_{|s|\to\infty}\frac{\|(is-A_{B})^{-1}\|}{|s|^{2\beta}}>0.$
    Now [11, Prop. 5.3] implies that $\limsup_{t\to\infty}t^{1/(2\beta)}\|T_{B}(t)\smash{A_{B}^{-1}}\|>0$, yielding the required contradiction.
@@
    Note that since $b\in L^{2}(0,1;\mathbb{R})$ we must have $(b_{n})_{n\geq 1}\in\ell^{2}$, so necessarily $\beta>1/2$ in Proposition 3.1. On the other hand, every $\beta>1/2$ can be achieved, for instance by considering the function $b$ whose Fourier sine series coefficients are $b_{n}=n^{-\beta}$ for $n\geq 1$. We refer the interested reader to [11, Rem. 6.4] for a discussion on the possibility of achieving these decay rates by means of functions $b$ that possess additional regularity.
@@ 3.2. The SCOLE model with non-linear damping
$2(1-a)\rho(x)-(\rho(x)\zeta(x))^{\prime}<-b,$
$E\!I(x)(1-a-2\zeta^{\prime}(x))+\frac{1}{2}(E\!I(x)\zeta(x))^{\prime}<-b$
@@
Consider the system (1.1) for the SCOLE model with non-linear damping as above. Suppose that $\phi\colon U\to U$ is monotone and locally Lipschitz continuous, and that (1.2) holds for some $\delta>0$. Then all generalised solutions of (1.1) satisfy $\|x(t)\|\to 0$ as $t\to\infty$, and all classical solutions satisfy $\|x(t)\|=O(t^{-1/2})$ as $t\to\infty$.
@@
By [20, Prop. 1.1] the operator AA is skew-adjoint, has compact resolvent, and its eigenvalues are simple. In addition, it follows from [20, Prop. 1.2] that the eigenvalues of AA are uniformly separated. Finally, $i\mathbb{R}\subseteq\rho(A_{B})$ and $\|(is-A_{B})^{-1}\|\lesssim 1+s^{2}$ for $s\in\mathbb{R}$ by [18, Thm. 3.1] and [9, Prop. 1.3], so result follows from Corollary 2.6. ∎

### v2

@@ 3.1. The wave equation with weak non-linear damping
to be solved subject to the boundary conditions $u(0,t)=u(1,t)=0$ for all $t>0$ and the initial conditions $u(\cdot,0)\in H_{0}^{1}(0,1)$, $u_{t}(\cdot,0)\in L^{2}(0,1)$. Here the function $b\in L^{2}(0,1;\mathbb{R})$ models the presence of weak (distributed) damping and the function $\phi\colon\mathbb{C}\to\mathbb{C}$ is potentially non-linear. We may formulate the problem in the form of (1.1) for the state variable $x(t)=(u(\cdot,t),u_{t}(\cdot,t))$, $t\geq 0$, by setting $X=H_{0}^{1}(0,1)\times L^{2}(0,1)$, $U=\mathbb{C}$, choosing $A$ to be the operator defined by $A(u,v)=(v,u^{\prime\prime})$ for all $(u,v)$ in the domain $D(A)=(H^{2}(0,1)\cap H_{0}^{1}(0,1))\times H_{0}^{1}(0,1)$ and defining $B\in\mathcal{B}(U,X)$ by $Bz=(0,b(\cdot)z)$ for all $z\in\mathbb{C}$. We denote the (rescaled) Fourier sine series coefficients of $b$ by
@@
Consider the system (1.1) for the weakly damped wave equation as above, and suppose there exists $\beta>0$ such that $|b_{n}|\gtrsim n^{-\beta}$ for all $n\geq 1$. Suppose further that $\phi\colon U\to U$ is monotone and locally Lipschitz continuous, and that (1.2) holds for some $\delta>0$. Then all mild solutions of (1.1) satisfy $\|x(t)\|\to 0$ as $t\to\infty$, and all classical solutions satisfy $\|x(t)\|=O(t^{-1/(2\beta)})$ as $t\to\infty$.
@@
The operator AA is skew-adjoint, and its spectrum consists of the simple eigenvalues $i\pi n$ for $n\in\mathbb{Z}\setminus\{0\}$, which are uniformly separated. As in the proof of [12, Cor. 6.3], our assumption on the decay of the Fourier sine series coefficients of $b$ implies that $i\mathbb{R}\subseteq\rho(A_{B})$ and $\|(is-A_{B})^{-1}\|\lesssim 1+|s|^{2\beta}$ for all $s\in\mathbb{R}$. The result now follows from Corollary 2.6. ∎
@@
    Note that in the particular case where $\phi$ is the identity map on $\mathbb{C}$, so that (1.2) holds for $\delta=1$, Proposition 3.1 is sharp in the sense that if $\limsup_{n\to\infty}n^{\beta}|b_{n}|>0$ then for any function $r\colon\mathbb{R}_{+}\to(0,\infty)$ such that $r(t)=o(t^{-1/(2\beta)})$ as $t\to\infty$ there exists a classical solution such that $\|x(t)\|\neq O(r(t))$ as $t\to\infty$. Indeed, if this is not the case then an application of the uniform boundedness principle gives $\|T_{B}(t)\smash{A_{B}^{-1}}\|=O(r(t))$ as $t\to\infty$. On the other hand, since $\limsup_{n\to\infty}n^{\beta}|b_{n}|>0$ it follows from [12, Prop. 5.1] that lim sup_—s—→∞∥(is-AB)-1∥—s—2β¿0. Now [12, Prop. 5.3] implies that $\limsup_{t\to\infty}t^{1/(2\beta)}\|T_{B}(t)\smash{A_{B}^{-1}}\|>0$, yielding the required contradiction.
@@
    Note that since $b\in L^{2}(0,1;\mathbb{R})$ we must have $(b_{n})_{n\geq 1}\in\ell^{2}$, so necessarily $\beta>1/2$ in Proposition 3.1. On the other hand, every $\beta>1/2$ can be achieved, for instance by considering the function $b$ whose Fourier sine series coefficients are $b_{n}=n^{-\beta}$ for $n\geq 1$. We refer the interested reader to [12, Rem. 6.4] for a discussion on the possibility of achieving these decay rates by means of functions $b$ that possess additional regularity.
@@ 3.2. The SCOLE model with non-linear damping
$2(1-a)\rho(x)-(\rho\zeta)^{\prime}(x)<-b,$
$E\!I(x)(1-a-2\zeta^{\prime}(x))+\frac{1}{2}(E\!I\,\zeta)^{\prime}(x)<-b$
@@
Consider the system (1.1) for the SCOLE model with non-linear damping as above. Suppose that $\phi\colon U\to U$ is monotone and locally Lipschitz continuous, and that (1.2) holds for some $\delta>0$. Then all mild solutions of (1.1) satisfy $\|x(t)\|\to 0$ as $t\to\infty$, and all classical solutions satisfy $\|x(t)\|=O(t^{-1/2})$ as $t\to\infty$.
@@
By [21, Prop. 1.1] the operator AA is skew-adjoint, has compact resolvent, and its eigenvalues are simple. In addition, it follows from [21, Prop. 1.2] that the eigenvalues of AA are uniformly separated. Finally, $i\mathbb{R}\subseteq\rho(A_{B})$ and $\|(is-A_{B})^{-1}\|\lesssim 1+s^{2}$ for $s\in\mathbb{R}$ by [19, Thm. 3.1] and [10, Prop. 1.3], so result follows from Corollary 2.6. ∎